OpenStudy (anonymous):

Verify using Sum-to-Product and Product-to-Sum Identities. 2cos^2 x/2 tanx=tanx+sinx

OpenStudy (anonymous):

\[2\cos ^2\frac{ x }{ 2 }\tan x =\tan x +\sin x\]

OpenStudy (anonymous):

I know to start with the difficult side first. LHS

OpenStudy (loser66):